OpenStudy (anonymous):

Find the limit. Use l'Hospital's Rule if appropriate. If there is a more elementary method, consider using it. lim (1+a/x)^bx x>infinite please i need help solving this and dont tell me do this step or this help me solve the problem all the way to the final answer

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to\infty}\left(1+\frac{a}{x}\right)^{bx}\]Hmm I think this relates to the definition of e. But let's just work through it and see what's going on.

zepdrix (zepdrix):

So right now we're getting the indeterminate form: \(\Large\rm 1^{\infty}\) yes?

zepdrix (zepdrix):

So there's a fancy little trick we can do. \[\Large\rm e^{\ln \color{orangered}{x}}=\color{orangered}{x}\]This is true since the exponential and log are inverse functions of one another. We want to use this but in reverse.\[\Large\rm \color{orangered}{\lim_{x\to\infty}\left(1+\frac{a}{x}\right)^{bx}}=e^{\ln\left[\color{orangered}{\lim_{x\to\infty}\left(1+\frac{a}{x}\right)^{bx}}\right]}\]

zepdrix (zepdrix):

So let's ignore the base for a few steps... and let's pass the limit out of the logarithm. So we're looking at this:\[\Large\rm \lim_{x\to\infty} \ln\left[\left(1+\frac{a}{x}\right)^{bx}\right]\] With me so far? Where you at broski?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Rules of logs allows us to bring the bx down in front,\[\Large\rm \lim_{x\to\infty} bx~ \ln\left(1+\frac{a}{x}\right)\]Let's take the b outside... And we're going to do another fancy thing, this time with our x. x is equal to the reciprocal of the reciprocal of x. \[\Large\rm x=\frac{1}{\left(\frac{1}{x}\right)}\]The reason we would want to do this is, it puts something in the denominator. We want a numerator and denominator so we can check out L'Hopital maybe.

zepdrix (zepdrix):

So here's what our stuff in the exponent looks like so far, \[\Large\rm b\lim_{x\to\infty} \frac{\ln\left(1+\frac{a}{x}\right)}{\frac{1}{x}}\]If we take the limit now, what form are we getting? (It should still be indeterminate, but what specific form?)

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

when i had put that final answer into webassign it didnt accept ln(1+a/x)/1/x

zepdrix (zepdrix):

correct. this simplifies down a lot further. I was asking you a question.

OpenStudy (anonymous):

oh okay

OpenStudy (akashdeepdeb):

Now apply L'hospital's rule. Because they both are in 0/0 form.